OpenStudy (marigirl):

can you please check over why my method is incorrect for this question. Topic differentiation A Cylinder of height h and radius r is inside a sphere of of radius 10. find the value of r that will maximise volume of cylinder

OpenStudy (marigirl):

Volume of a cylinder is v=pi r^2*h so i went about getting an expression or h to substitute in by using similar triangles

OpenStudy (marigirl):

so i used similar triangles and obtained h=20-2r .. the model answers showed that they used pythagoras.. why is my methods incorrect?

OpenStudy (amistre64):

height is based off a curve, not a line

OpenStudy (amistre64):

height = sqrt(10^2 - x^2)

OpenStudy (amistre64):

oh, and even tho thats pythagorean, its really based off of the equation for the circle of the sphere

OpenStudy (marigirl):

but if u take the into consideration the height and radius measurements from the origin of x and y axis - wont it be a straight line?

OpenStudy (amistre64):



OpenStudy (amistre64):

all points equidistant from the center ... in this case the origin ... is the equation x^2 + y^2 = r^2

OpenStudy (amistre64):

height is measured as 2y in this case. radius is x

OpenStudy (marigirl):

but if u take the into consideration the height and radius measurements from the origin of x and y axis - wont it be a straight line?

OpenStudy (amistre64):

repeating faulty arguments do not make them any more correct ...

OpenStudy (loser66):